OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own. @MrNood

OpenStudy (anonymous):

Let x be how long it takes Jennifer to do job, then 1/x+ (1/(x-3))=1/2, So x=6, So it would take Jennifer 6 hours.

OpenStudy (anonymous):

but how would i show work im lost

OpenStudy (mrnood):

as usual @amonaroll32 too quick to show that hey can answer - and no help to the OP

OpenStudy (anonymous):

will you help me show work? @amonaroll32 @MrNood

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

Hey I gave you an answer look at your closed post

OpenStudy (anonymous):

Sorry I didn't realize your post was closed

OpenStudy (mrnood):

OK - the amount of work to be done is "1 room" so if Jennifer takes x hours to do th ework then her rate of work is 1room/xhours OK so far?? and Carl takes x-3 hours so HIS rate of work is 1 room/x-3

OpenStudy (anonymous):

how do i make it not closed i just started using open study @desiswag

OpenStudy (anonymous):

I don't think you can make it open again, but you can still view the post

OpenStudy (anonymous):

@MrNood okay i got 1/x +1/x-3=1/2 so far what would i do next

OpenStudy (mrnood):

Now - we are told when they work together that the room is done in 2 hours so the total rate of work is 1room/2 hours so the total work is 1/x+1/(x-3) = 1/2

OpenStudy (mrnood):

\[\frac{ 1 }{ x }+ \frac{ 1 }{ (x-3) } = \frac{ 1 }{ 2}\] yo can solve that for x

OpenStudy (anonymous):

how would i do that though? @mrnood

OpenStudy (anonymous):

@MrNood

OpenStudy (mrnood):

I am not about to give oyu the complete solution. This is a simple equation involving common denominators and re-arranging to solve for x I guess that is the subject you are currently studying - so give it a try for yourself (hint - first put the left side over a common denominator....)

OpenStudy (anonymous):

alright so far i got 2x+6/2x^2+6x + 2x/2x^2+6x=x^2+3x/2x^2+6x

OpenStudy (anonymous):

wait dont you times both sides by 2 i think i did it wrong @MrNood

OpenStudy (mrnood):

Take it steady - FIRST sort out the common denominator

OpenStudy (anonymous):

i dont g=know what to put for the common denominator

OpenStudy (anonymous):

get***

OpenStudy (mrnood):

I'm not sure I'm up for explaining how to solve equations. start here: \[\frac{ a }{ b} + \frac{ c}{ d } =\frac{ =(ad + cb) }{ bd }\] the easiest common denominator is got by multiplying the two existing denominators